a battery of emf 15V and internal resistant 3ohm is connected to two resistors 3ohm and 6  ohm connected in parallel       find p.d. between the terminals of the battery

Grade:10

1 Answers

DR STRANGE
7 years ago
HI Reema,
total resistance R = 3 + (3*6)/(3+6) = 5 ohm
emf = 15V 
current though terminal of battery i = V/R = 15/5 =3A 
p.d. between terminals = emf – i*(internal resistance) = 15 – 3*3 =15 – 9 = 6V
Approve if you are satisfied :)